OpenStudy (anonymous):

What is the solution of the equation? 4.7x + 3.8 = 13.2 A. 3.6 B. 3.7 C. 2 D. 3 need help i want to show my son how to do it but i failed in math so can someone run me through this so i can show him thanks

OpenStudy (cwrw238):

4.7x + 3.8 = 13.2 first aim is to get 4.7x on left of the '=' and numbers on right to do this subtract 3.8 from both sides: 4.7x + 3.8 - 3.8 = 13.2 - 3.8 4.7x = 9.4 now divide both sides by 4.7 to get value of x: x = 9.4/4.7

OpenStudy (cwrw238):

so divide 9.4 by 4.7 and you have x

OpenStudy (anonymous):

2

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

ty

OpenStudy (cwrw238):

yw